# Fisher-Rao Gradient Flow: Geodesic Convexity and Functional Inequalities

### v1

@@ 1.3. Main results
Firstly, we show that the gradient dominance condition and the geodesic convexity do not hold for the Fisher-Rao gradient flows of the KL divergence (1.16). This poses significant challenges for establishing the convergence of the dynamics using the functional inequality approach. The informal statement is presented below and the detailed analysis is in Theorems 3.3 and 4.1.
@@
Secondly, for general $f$-divergences as the energy functionals, we derive a sufficient condition for geodesic convexity under the Fisher-Rao geometry, and a sufficient and necessary condition for the gradient dominance condition to hold. These lead to general convergence results for the corresponding Fisher-Rao gradient flows. We state the informal result below and the detailed result is in Theorems 3.4, 4.2 and 4.6; our proofs use novel techniques from linear programming in optimization to reduce the challenging global nonlocal interaction to a three-point interaction which is much simpler to deal with.
@@
Under the Fisher-Rao geometry, the $f$-divergence is geodesically convex when $xf^{\prime}(x)$ is concave. Additionally under the condition $f^{\prime\prime}(1)>0$, the $f$-divergence achieves $\alpha_{f}$-strong geodesic convexity with $\alpha_{f}>0$ depending solely on $f$.
Moreover, for general $f$-divergences, the gradient dominance condition holds with an $\alpha_{f}>0$ depending solely on $f$ if and only if
$\begin{split}\exists\alpha>0\qquad\big{(}f^{\prime}(y)-f^{\prime}(x)\bigr{)}^{2}\geq\alpha\left(\frac{1}{x}-\frac{1}{y}\right)\left(\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\right)\quad\text{for all }\ 0<x<1<y.\end{split}$
An sufficient condition is that there exists a constant $\alpha_{s}>0$ such that $x^{2}f^{\prime\prime}(x)>\alpha_{s}$ for $0<x\leq 1$. Note that the KL divergence does not satisfy the conditions.
Thirdly, we study the case of the KL divergence in further depth and seek other possible functional inequalities to establish the convergence. We prove a novel dual type of gradient dominance conditions that allow us to obtain the convergence of the Fisher-Rao gradient flows for general $f$-divergences. To the best of our knowledge, such dual type of functional inequalities is not known in the Wasserstein case and it is a nontrivial, unique discovery in the Fisher-Rao case. The informal statement is as follows; the detailed statement is in Section 5.
@@
(1.20) $\frac{\rm d}{{\rm d}t}D_{\bar{f}}[\rho_{t}\|{\rho^{*}}]\leq-\alpha_{f}\Bigl{(}D_{\bar{f}}[\rho_{t}\|{\rho^{*}}]+D_{f}[\rho_{t}\|{\rho^{*}}]\Bigr{)}.$
@@
For $f\in C^{2}(0,+\infty)$ and $f^{\prime\prime}(1)>0$, the above inequality holds with an $\alpha_{f}$ independent of $\rho^{*}$, if we set ${\bar{f}}(x)=\frac{(x-1)^{2}}{x}$.
@@
Furthermore, for the KL divergence, if we take ${\bar{f}}(x)=xf(\frac{1}{x})$, then the dual gradient dominance condition holds with an explicit $\alpha_{f}=1$. This can also be generalized to $f$-divergences with $f$ being a polynomial.
@@
With these established functional inequalities, it is straightforward to show the corresponding gradient flows converge to the target exponentially fast at the same rate, given that the gradient flows are well posed, allowing us to differentiate the energy functional along the dynamics. The main focus of this paper is the rigorous proof of the functional inequalities and we leave the study of the well-posedness of the gradient flows for future work.
1.4. Related works
1.4.1. Displacement convexity and functional inequalities
@@
With respect to the Wasserstein metric, geodesic convexity is equivalent to the displacement convexity, a concept introduced by McCann in [64]. Further exploration of displacement convexity on Riemannian manifolds and more general spaces can be found in [69, 25, 70, 82, 58, 66]. Additionally, the work [16] studies the displacement convexity concerning the generalized Wasserstein metric with nonlinear mobility. These concepts have been extended to more general settings such as graphs, discrete space, Markov chains with other metrics, see for instance [31, 32, 30].
@@
Various functional inequalities related to the Wasserstein metric have been developed, such as the Poincaré inequality, log-Sobolev inequality, and Talagrand’s inequality; see [72, 80, 69]. These functional inequalities can be used to study the convergence of Wasserstein gradient flows and related PDEs [69]. Log-Sobolev type inequalities have been proved for general divergences or relative entropy and for general nonlinear and nonlocal McKean-Vlasov equations; see for instance [5, 15, 17, 3]. However, few functional inequalities have been developed for the Fisher-Rao gradient flow.
1.4.2. Fisher-Rao gradient flow
The Fisher-Rao metric and its finite-dimensional counterpart in parameter space, the Fisher information matrix, are fundamental concepts in information geometry [2, 7]; see also [74, 35, 79]. The geodesic distance under the Fisher-Rao metric is known as the Fisher-Rao distance, which, under a mass conservation constraint, is also referred to as the spherical Hellinger distance [53, 52, 12]. The dynamical formulation of the Fisher-Rao distance is related to the development of unbalanced optimal transport and the Wasserstein-Fisher-Rao distance [24, 51, 54].
The Fisher-Rao gradient flow of the KL divergence (1.16) is a mean-field equation of birth-death type [59, 60]:
$\frac{\partial\rho_{t}}{\partial t}=-\rho_{t}\Lambda_{t},\qquad\Lambda_{t}=\bigl{(}\log\rho_{t}-\log{\rho^{*}}\bigr{)}-\mathbb{E}_{\rho_{t}}[\log\rho_{t}-\log{\rho^{*}}],$
@@ 1.4.2. Fisher-Rao gradient flow
where the birth-death rate $\Lambda_{t}$ contains a nonlocal interaction term that allows global movement of mass. Similar equations in discrete space, known as the replicator equations, have been studied in evolutionary game theory [43]. Projection of the Fisher-Rao gradient flows to a parametric family of distributions leads to natural gradient flows. Time discretizations of these flows, known as natural gradient descent algorithms, have been widely used in variational inference [1, 55, 62, 21].
@@
Fisher-Rao gradient flows have found applications in various fields. For example, the corresponding birth-death process has been used in sequential Monte Carlo samplers to reduce the variance of particle weights [27], and in the acceleration of Langevin sampling [59, 60] and training of neural networks [75]. The Fisher-Rao gradient flow of the $\chi^{2}$ distance arises as a mean-field equation for ensemble MCMC methods that address multimodal distributions [56]. Kernel approximations of Fisher-Rao gradient flows have been proposed for sampling [63]. Furthermore, parametric approximation of Fisher-Rao gradient flows employing Kalman’s methodology [49, 33, 48] has led to the development of efficient, derivative-free posterior approximation algorithms. These algorithms, based on Gaussian or Gaussian mixture approximations [44, 22], have proven particularly useful for Bayesian inverse problems. Other applications in filtering [40] and reinforcement learning [50] have also been explored in the literature.
1.5. Organization of this paper
@@
2. Fisher-Rao Gradient Flows of $f$-divergences
@@
(2.1) $\mathcal{E}(\rho)=D_{f}[\rho\|{\rho^{*}}]=\int{\rho^{*}}f\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\,\mathrm{d}\theta,$
@@
(2.2) $D_{f}[\rho\|{\rho^{*}}]\geq\int{\rho^{*}}\Bigl{(}f(1)+f^{\prime}(1)(\rho/{\rho^{*}}-1)\Bigr{)}\,\mathrm{d}\theta\geq-\int\bigl{|}f^{\prime}(1)(\rho-{\rho^{*}})\bigr{|}\,\mathrm{d}\theta\geq-2\lvert f^{\prime}(1)\rvert.$
This means $f$-divergence is always well defined (with a possible value +$\infty$+ ∞). Moreover, Jensen’s inequality implies that ${\rho^{*}}$ is the unique minimizer. The KL-divergence is a special $f$-divergence with $f=x\log x$.
For the $f$-divergence, its first variation with respect to $\rho$ is $f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}-\mathbb{E}_{\rho}[f^{\prime}(\frac{\rho}{{\rho^{*}}})]$ where we subtract a constant so that its mean is zero. Using the formula of the Fisher-Rao metric (1.15), we get the corresponding Fisher-Rao gradient flow as
(2.3) $\frac{\partial\rho_{t}}{\partial t}=-M(\rho_{t})^{-1}\frac{\delta\mathcal{E}}{\delta\rho}\Bigr{|}_{\rho_{t}}=-\rho_{t}f^{\prime}\Bigl{(}\frac{\rho_{t}}{{\rho^{*}}}\Bigr{)}+\rho_{t}\mathbb{E}_{\rho_{t}}\Bigl{[}f^{\prime}\Bigl{(}\frac{\rho_{t}}{{\rho^{*}}}\Bigr{)}\Bigr{]}.$
We note that the Fisher-Rao metric (1.15) stands out as the unique metric, up to constants, that satisfies the diffeomorphism invariance property [19, 6, 10]. Under the Fisher-Rao metric, these gradient flows are expected to maintain this invariance. The diffeomorphism invariance implies that the convergence property of Fisher-Rao gradient flows for a general target density ${\rho^{*}}$ aligns with that for the standard Gaussian target density, and may achieve a convergence rate independent of ${\rho^{*}}$. For this reason, they have been widely used in sampling applications.
@@
(2.4) $\frac{\partial\rho_{t}}{\partial t}=-\rho_{t}\Bigl{(}\log\rho_{t}-\log{\rho^{*}}-\mathbb{E}_{\rho_{t}}[\log\rho_{t}-\log{\rho^{*}}]\Bigr{)},$
which is the most commonly used Fisher-Rao gradient flow for sampling applications [59, 21, 60, 28, 85]. When $f(x)=(x-1)^{2}$ which corresponds to the $\chi^{2}$ divergence, we get
(2.5) $\frac{\partial\rho_{t}}{\partial t}=-\rho_{t}\Bigl{(}\frac{\rho_{t}}{\rho^{*}}-\mathbb{E}_{\rho_{t}}[\frac{\rho_{t}}{\rho^{*}}]\Bigr{)},$
which has been used in the context of ensemble samplers [56].
3. Geodesic Convexity in the Fisher-Rao Geometry
In this section, we explore the geodesic convexity of $f$-divergences in the Fisher-Rao geometry. We will present counterexamples to demonstrate that the KL divergence lacks geodesic convexity. Subsequently, we will provide sufficient conditions to identify the specific $f$-divergences that achieve geodesic convexity. The discussions in this section parallel those of displacement convexity in the Wasserstein geometry. The geometric calculations are primarily carried out in a formal basis  [^1] However, we note that they can be made rigorous by restricting to probability simplex $\sum_{i=1}^{K}\rho_{i}=1,\rho_{i}>0$, where the Riemannian geometry is rigorously defined.. We first outline the preliminary concepts in the Euclidean space and abstract manifolds and then move to the Fisher-Rao geometry.
3.1. Preliminaries on convex analysis

### v2

@@ 1.3 Main results
Firstly, we prove that the gradient dominance condition and the geodesic convexity do not hold for the Fisher-Rao gradient flows of the KL divergence (16). This poses significant challenges for establishing the convergence of the dynamics using the functional inequality approach. The informal statement is presented below and the detailed analysis is in Theorems 3.1 and 4.1. Related results have been studied in the concurrent work [ZhuMielke2024, Lemma 6.2] which focuses on the setting without the mass preservation constraint.
@@
Secondly, for general $f$-divergences as the energy functionals, we derive a sufficient condition on $f$ for geodesic convexity under the Fisher-Rao geometry, and a sufficient and necessary condition (21) on $f$ for the gradient dominance (see details in 4.1) to hold. These lead to general convergence results for the corresponding Fisher-Rao gradient flows. We state our informal result below and the detailed result is in Theorems 3.3, 4.3, and 4.8. We note that in contemporary work, [mielke2025hellinger, Section 4] studies the gradient dominance condition for a particular class of $f$-divergences with $f^{\prime\prime}(x)=x^{p}$.
@@
Under the Fisher-Rao geometry, the $f$-divergence is geodesically convex for any $\rho^{*}$} when $xf^{\prime}(x)$ is concave. Additionally under the condition $f^{\prime\prime}(1)>0$, the $f$-divergence achieves $\alpha_{f}$-strong geodesic convexity for any $\rho^{*}$} with $\alpha_{f}>0$ depending solely on $f$.
Moreover, for general $f$-divergences, the gradient dominance condition holds for any $\rho^{*}$} with an $\alpha_{f}>0$ depending solely on $f$ if and only if
(21) $\begin{split}\exists\alpha>0\qquad\Bigl(f^{\prime}(y)-f^{\prime}(x)\Bigr)^{2}\geq\alpha\left(\frac{1}{x}-\frac{1}{y}\right)\left(\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\right)\quad\end{split}$
for all $0<x<1<y$. An sufficient condition is that there exists a constant $\alpha_{s}>0$ such that $x^{2}f^{\prime\prime}(x)>\alpha_{s}$ for $0<x\leq 1$. Note that the KL divergence does not satisfy the conditions.
The form of the sufficient and necessary condition may seem hard to parse at first glance. It can be derived from a special case of gradient dominance when $\rho$ and $\rho^{*}$} are two-point discrete densities; for more details see the proof for Theorem 4.3.
Thirdly, we study the case of the KL divergence in further depth and seek other possible functional inequalities to establish the convergence, thereby extending previous results. We prove a novel dual type of gradient dominance conditions that allow us to obtain the convergence of the Fisher-Rao gradient flows for general $f$-divergences. To the best of our knowledge, such dual type of functional inequalities is not known in the Wasserstein case and it is a new property found in the Fisher-Rao case.
The informal statement is as follows; the detailed statement is in Section 5. Related work [yao2024minimizing, Theorem 1] establishes that the KL divergence is a relatively strongly convex functional (in the sense of [lu2018relatively]) with respect to the KL divergence itself, which leads to the convergence of the gradient flow under the reverse KL divergence, a case that can be covered by our result; see details in Theorem 5.5.
@@
(22) $\frac{\rm d}{{\rm d}t}D_{\bar{f}}[\rho_{t}\|{\rho^{*}}]\leq-\alpha_{f}\Bigl(D_{\bar{f}}[\rho_{t}\|{\rho^{*}}]+D_{f}[\rho_{t}\|{\rho^{*}}]\Bigr).$
@@
For $f\in C^{2}(0,+\infty)$ and $f^{\prime\prime}(1)>0$, the above inequality holds with a constant $\alpha_{f}$ depending only on $f$ and independent of $\rho$ and $\rho^{*}$, if we take ${\bar{f}}(x)=\frac{(x-1)^{2}}{x}$. In this case, the dual divergence $D_{{\bar{f}}}$ corresponds to the reverse $\chi^{2}$ divergence.
@@
Furthermore, for the KL divergence, if we take ${\bar{f}}(x)=xf(\frac{1}{x})$, which corresponds to the reverse KL divergence, the dual gradient dominance condition holds with an explicit $\alpha_{f}=1$. This can also be generalized to $f$-divergences with $f$ being a polynomial.
@@
With these established functional inequalities, it is straightforward to show the corresponding gradient flows converge to the target exponentially fast at the same rate, given that the gradient flows are well posed, allowing us to differentiate the energy functional along the dynamics. The main focus of this paper is the rigorous proof of the functional inequalities under appropriate sufficient and/or necessary conditions, while we leave the study of the well-posedness of the gradient flows for future work.
1.4 Related works
1.4.1 Displacement convexity and functional inequalities
@@
With respect to the Wasserstein metric, geodesic convexity is equivalent to the displacement convexity, a concept introduced by McCann in [mccann1997convexity]. Further exploration of displacement convexity on Riemannian manifolds and more general spaces can be found in [otto2000generalization, cordero2001riemannian, otto2005eulerian, von2005transport, lott2009ricci, ohta2011displacement]. Additionally, the work [carrillo2010nonlinear] studies the displacement convexity concerning the generalized Wasserstein metric with nonlinear mobility. These concepts have been extended to more general settings such as graphs, discrete space, Markov chains with other metrics, see for instance [EM14, EMM19, EFMM22].
@@
Various functional inequalities related to the Wasserstein metric have been developed, such as the Poincaré inequality, log-Sobolev inequality, and Talagrand’s inequality; see [poincare1890equations, talagrand1996transportation, otto2000generalization]. These functional inequalities can be used to study the convergence of Wasserstein gradient flows and related PDEs [otto2000generalization]. Log-Sobolev type inequalities have been proved for general divergences or relative entropy and for general nonlinear and nonlocal McKean-Vlasov equations; see for instance [AMTU, carrillo2001entropy, CMV03, ambrosio2005gradient]. However, few functional inequalities have been developed for the Fisher-Rao gradient flow.
1.4.2 Fisher-Rao gradient flow
The Fisher-Rao metric is widely used in finite-dimensional parametric density spaces, where it reduces to the Fisher information matrix, a fundamental concept in information geometry [amari2016information, ay2017information]; see also [rao1945information, FRmetricInfDim1991, srivastava2007riemannian]. In probability density spaces, Fisher-Rao and Wasserstein metrics have been studied in combination, leading to the Wasserstein-Fisher-Rao (also known as Hellinger–Kantorovich) metric [liero2018optimal, kondratyev2016new, chizat2018interpolating, chizat2018unbalanced]. A key advantage of this framework is its ability to handle unnormalized probability densities that are not constrained to integrate to one, giving rise to unbalanced optimal transport, which has found applications across various domains. The interpolation can also be defined via a generalization of the dynamical Benamou-Brenier formulation [benamou2000computational] by incorporating a source term measured by the Fisher-Rao metric in the continuity equation. The associated gradient flows [kondratyev2019spherical] and functional inequalities [kondratyev2020convex, mielke2025hellinger] have been studied following Otto’s formalism. In the present work, we focus on the Fisher–Rao metric for normalized probability densities and the associated gradient flows preserve the probability mass.
The Fisher-Rao gradient flow of the KL divergence (16) takes the form
$\frac{\partial\rho_{t}}{\partial t}=-\rho_{t}\Lambda_{t},\qquad\Lambda_{t}=\bigl(\log\rho_{t}-\log{\rho^{*}}\bigr)-\mathbb{E}_{\rho_{t}}[\log\rho_{t}-\log{\rho^{*}}].$
@@ 1.4.2 Fisher-Rao gradient flow
It is a mean-field equation of birth-death type [lu2019accelerating, lu2022birth] where the birth-death rate $\Lambda_{t}$ contains a nonlocal interaction term that allows global movement of mass. Similar equations in discrete space, known as the replicator equations, have been studied in evolutionary game theory [hofbauer2003evolutionary]. Projection of the Fisher-Rao gradient flows to a parametric family of distributions leads to natural gradient flows. Time discretizations of these flows, known as natural gradient descent algorithms, have been widely used in variational inference [amari1998natural, lin2019fast, martens2020new, chen2023sampling].
@@
Fisher-Rao gradient flows have found applications in various fields. For example, the corresponding birth-death process has been used in sequential Monte Carlo samplers to reduce the variance of particle weights [del2006sequential], and in the acceleration of Langevin sampling [lu2019accelerating, lu2022birth] and training of neural networks [rotskoff2019global]. The Fisher-Rao gradient flow of the $\chi^{2}$ divergence arises as a mean-field equation for ensemble MCMC methods that address multimodal distributions [lindsey2022ensemble]. Kernel approximations of Fisher-Rao gradient flows have been proposed for sampling [maurais2024sampling, zhu2024kernel]. Furthermore, parametric approximation of Fisher-Rao gradient flows employing Kalman’s methodology [kalman1960new, evensen1994sequential, julier1997new] has led to the development of efficient, derivative-free posterior approximation algorithms. These algorithms, based on Gaussian or Gaussian mixture approximations [huang2022efficient, chen2024efficient, che2025stable], have proven particularly useful for Bayesian inverse problems. Other applications in filtering [halder2018gradient] and reinforcement learning [kerimkulov2023fisher] have also been explored in the literature.
1.5 Organization of this paper
@@
2 Fisher-Rao Gradient Flows of $f$-divergences
@@
(23) $\mathcal{E}(\rho)=D_{f}[\rho\|{\rho^{*}}]=\int{\rho^{*}}f\Bigl(\frac{\rho}{{\rho^{*}}}\Bigr)\,\mathrm{d}\theta,$
@@
(24) $D_{f}[\rho\|{\rho^{*}}]\geq\int{\rho^{*}}\Bigl(f(1)+f^{\prime}(1)(\rho/{\rho^{*}}-1)\Bigr)\,\mathrm{d}\theta\geq-\int\bigl|f^{\prime}(1)(\rho-{\rho^{*}})\bigr|\,\mathrm{d}\theta\geq-2\lvert f^{\prime}(1)\rvert$.
This means $f$-divergence is always well defined (with a possible value $\infty$). Moreover, Jensen’s inequality implies that $\rho^{*}$} is the unique minimizer. The KL-divergence is a special $f$-divergence with $f=x\log x$.
For the $f$-divergence, its first variation with respect to $\rho$ is $f^{\prime}\bigl(\frac{\rho}{{\rho^{*}}}\bigr)-\mathbb{E}_{\rho}[f^{\prime}(\frac{\rho}{{\rho^{*}}})]$ where we subtract a constant so that its mean is zero. Using the formula of the Fisher-Rao metric (15), we get the corresponding Fisher-Rao gradient flow as
(25) $\frac{\partial\rho_{t}}{\partial t}=-M(\rho_{t})^{-1}\frac{\delta\mathcal{E}}{\delta\rho}\Bigr|_{\rho_{t}}=-\rho_{t}f^{\prime}\Bigl(\frac{\rho_{t}}{{\rho^{*}}}\Bigr)+\rho_{t}\mathbb{E}_{\rho_{t}}\Bigl[f^{\prime}\Bigl(\frac{\rho_{t}}{{\rho^{*}}}\Bigr)\Bigr].$
We note that the Fisher-Rao metric (15) stands out as the unique metric, up to constants, that satisfies the diffeomorphism invariance property [cencov2000statistical, ay2015information, bauer2016uniqueness]. Under the Fisher-Rao metric, these gradient flows are expected to maintain this invariance. The diffeomorphism invariance implies that the convergence property of Fisher-Rao gradient flows for a general target density $\rho^{*}$} aligns with that for the standard Gaussian target density, and may achieve a convergence rate independent of $\rho^{*}$}.
@@
(26) $\frac{\partial\rho_{t}}{\partial t}=-\rho_{t}\Bigl(\log\rho_{t}-\log{\rho^{*}}-\mathbb{E}_{\rho_{t}}[\log\rho_{t}-\log{\rho^{*}}]\Bigr),$
which is the most commonly used Fisher-Rao gradient flow for sampling applications [lu2019accelerating, chen2023sampling, lu2022birth, domingo2023explicit, yan2023learning]. When $f(x)=(x-1)^{2}$ which corresponds to the $\chi^{2}$ divergence, we get
(27) $\frac{\partial\rho_{t}}{\partial t}=-\rho_{t}\Bigl(\frac{\rho_{t}}{\rho^{*}}-\mathbb{E}_{\rho_{t}}[\frac{\rho_{t}}{\rho^{*}}]\Bigr),$
which has been used in the context of ensemble samplers [lindsey2022ensemble].
Note that in (27), evaluating the right-hand side appears to require knowledge of the normalization constant of $\rho^{*}$, which is typically unavailable in sampling applications. In fact, it has been demonstrated in [chen2023sampling] that the KL divergence is unique among $f$-divergences, up to constants, in that gradient flows resulting from it do not depend on the normalization constant of the target distribution. Nevertheless, in (27), the normalization constant appears as a multiplicative factor, and we can absorb it into the time variable and rescale the dynamics so that the right-hand side no longer depends on this constant; this approach changes the time scale. In [lindsey2022ensemble], ensemble approximation is employed with birth-death type updating rules to a slight variant of (27), which does not require the normalization constant to implement.
3 Geodesic Convexity in the Fisher-Rao Geometry
In this section, we explore the geodesic convexity of $f$-divergences in the Fisher-Rao geometry. We will present counterexamples to demonstrate that the KL divergence lacks geodesic convexity. Subsequently, we will provide sufficient conditions to identify the specific $f$-divergences that achieve geodesic convexity. The discussions in this section parallel those of displacement convexity in the Wasserstein geometry. The geometric calculations are primarily carried out formally, assuming the well-posedness of the geodesics and the gradient flow. We first outline the preliminary concepts in the Euclidean space and abstract manifolds and then move to the Fisher-Rao geometry.
3.1 Preliminaries on convex analysis